X to the third power -3x equal 45

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

X to the third power -3x equal 45
------
x^3 - 3x = 45
----
x^3 - 3x - 45 = 0
-------
One Real zero at x = 3.8375
===============
Cheers,
Stan H.

Answer by Alan3354(69443)

X to the third power -3x equal 45
------------
Use ^ (Shift 6) for exponents.
x^3 - 3x = 45
-----------------
x =~ +8.37386354243376
x =~ -5.37386354243376
Using graphs
-------------
The 3rd root is complex